If the events A and B have P ( A ) = .7, P ( B ) = .4, and P ( A | B ) = .5. What is P ( A ∪ B )?
Answer #1

Answer:-

Given that:-

P(A)=0.7 ,P(B)=0.4 and P(A/B)=0.5 What P(A\cup B)?

P(A\cap B)=P(A/B)P(B)

P(A\cap B)=(0.5)(0.4)

P(A\cap B)=0.2

P(A\cup B)=P(A)+P(B)-P(A\cap B)

=0.7+0.4-0.2

P(A\cup B)=0.9